Solve 12(3+y)=5(2y+8)

Answer by macston(5194)

12(3+y)=5(2y+8)
36+12y=10y+40 Subtract 36 from each side.
12y=10y+4 Subtract 10y from each side.
2y=4 Divide each side by 2.
y=2
CHECK:
12(3+y)=5(2y+8)
12(3+2)=5(2(2)+8)
12(5)=5(4+8)
60=60